If applied potential across X ray tube is made 2/5 times then the minimum wavelength of X ray shifted by 1A.find original values of minimum Wavelength and applied potential?

Grade:12th pass

1 Answers

Ravikumar
4 years ago
V/2÷5(V)=lamda+1/lamda; 5/2=lamda+1/lamda;5 lamda=2lamda+2;lamda=0.66 which mentioned as minimum wavelength given.... Stopping potential =12400/0.66=18kev...U got it!!!!👍👍👍all the best for ur competitive exam ur preparing